# Zhang, Liu, and Zhao

### v3

@@
Focusing on stochastic programming (SP) with covariate information, this paper proposes an empirical risk minimization (ERM) method embedded within a nonconvex piecewise affine decision rule (PADR), which aims to learn the direct mapping from features to optimal decisions. We establish the nonasymptotic consistency result of our PADR-based ERM model for unconstrained problems and asymptotic consistency result for constrained ones. To solve the nonconvex and nondifferentiable ERM problem, we develop an enhanced stochastic majorization-minimization algorithm and establish the asymptotic convergence to (composite strong) directional stationarity along with complexity analysis. We show that the proposed PADR-based ERM method applies to a broad class of nonconvex SP problems with theoretical consistency guarantees and computational tractability. Our numerical study demonstrates the superior performance of PADR-based ERM methods compared to state-of-the-art approaches under various settings, with significantly lower costs, less computation time, and robustness to feature dimensions and nonlinearity of the underlying dependency.
@@ 1 Introduction
Pioneered by George Dantzig in the mid-1950s, stochastic programming (SP) provides a class of optimization models for decision-making under uncertainty by treating the uncertainty as a random variable whose probability distribution is usually unknown in practice. With the growing real-world applications and availability of large datasets, the field of SP has flourished with many theoretical and algorithmic advances (Birge and Louveaux 2011, Shapiro et al. 2021), in which the empirical probability distribution based on historical scenarios plays a central role in approximating the true distribution. However, the empirical distribution may provide poor estimation when the random variable is related to covariate information (also referred to as contextual information or features) which can be observed before the decision is made. An unconstrained SP with the covariate information $X=x\in\mathbb{R}^{p}$ is formulated as follows,
$\inf_{z\in\mathbb{R}^{d}}\quad\mathbb{E}_{Y}\left[\varphi({z};{Y})\mid X=x\right],$ (1)
where $z\in\mathbb{R}^{d}$ is a decision vector, $\varphi:\mathbb{R}^{d}\times\mathcal{Y}\to\mathbb{R}$ is a cost function, and $Y$ is a random variable supported by $\mathcal{Y}$ whose distribution is related to a feature vector $X$ supported by $\mathcal{X}\subseteq\mathbb{R}^{p}$. A simple instance is an inventory planning problem, in which $z$ represents order decisions for $d$ products, $\varphi$ represents the inventory cost, $Y$ represents uncertain demands, and $X$ represents covariate information such as seasonality, promotion, and weather. With historical scenarios of $(X,Y)$, a classical sample average approximation (SAA) model disregards the dependency of $Y$ on $X$ and thus may lead to inconsistent decisions; such deficiency is also observed in numerical results of Ban and Rudin (2018) and Bertsimas and Koduri (2021). Motivated by the latent dependency of the optimal solution to (1) on the feature vector, this paper aims to construct a data-driven decision rule that directly maps features to optimal decisions so as to approximately solve the SP problem (1) for any possible observation of the feature.
@@ 1.1 Problem Setting
$\inf_{f\in\mathcal{F}}\quad R(f)\coloneqq\mathbb{E}_{X,Y}\left[\varphi({f(X)};{Y})\right],$ (2)
@@
where $\mathcal{F}$ contains all measurable functions mapping from the feature space $\mathcal{X}$ to the decision space $\mathbb{R}^{d}$. By the interchangeability principle (Shapiro et al. 2021, Theorem 7.80), if a mapping $f^{*}\in\mathcal{F}$ is optimal to problem (2), then the decision $f^{*}(x)$ is optimal to (1) for almost every $x\in\mathcal{X}$. In general, problem (2) is intractable due to the infinite-dimensional function space and the unknown joint distribution of $X$ and $Y$. A common approximation method is to construct an empirical risk minimization (ERM) problem based on a dataset $S\coloneqq\{(x^{s},y^{s})\}_{s=1}^{n}$ with the decision rule $f$ restricted to a parametric DR class $\mathcal{H}$ (also known as hypothesis class):
$\inf_{f\in\mathcal{H}}\quad\widehat{R}_{S}(f)\coloneqq\frac{1}{n}\sum_{s=1}^{n}\varphi({f(x^{s})};{y^{s}}).$ (3)
@@
In this paper, we adopt the above DR-based ERM framework, recognizing that the efficacy of such a framework is influenced by the trade-off between approximation accuracy of $\mathcal{H}$ to the optimal DR $f^{*}$ and computational complexity of solving (3). For example, problem (3) with a convex cost function can be solved efficiently with $\mathcal{H}$ being the class of linear decision rules (LDR) (Beutel and Minner 2012, Ban and Rudin 2018), while the LDR has difficulty describing the potential nonlinear relation between decisions and features. On the other side, a DR class consisting of neural networks could have a strong approximation capability (Qi et al. 2022); nevertheless, the resulting nonconvex problem (3) with a multi-layer composite structure is difficult to solve with theoretical optimality guarantees, and the model may face overfitting issues when the dataset is not large enough.
@@
Therefore, this paper aims to address a critical question: what could be an amenable hypothesis class for the DR-based ERM framework so as to maintain both high reliability and computational efficiency for solving SP with covariate information? To this end, we employ in the ERM problem (3) a piecewise affine decision rule (PADR) that is broader than LDR but simpler than general nonlinear DRs. According to Scholtes (2012), any continuous piecewise affine (PA) function $f:\mathcal{X}\rightarrow\mathbb{R}$ can be expressed as the difference of max-affine functions:
$f(x;\theta)=\max_{k\in[K]}\left\{(\alpha^{k})^{\top}x+a^{k}\right\}-\max_{k\in[K]}\left\{(\beta^{k})^{\top}x+b^{k}\right\},$ (4)
@@
where $\theta\coloneqq\left(\alpha^{k},a^{k},\beta^{k},b^{k}\right)_{k=1}^{K}\in\mathbb{R}^{2K(p+1)}$ and $[K]\coloneqq\{1,\dots,K\}$. Throughout the theoretical analysis in this paper, we assume that the numbers of pieces in the two max-affine components of (4) are equal to simplify the exposition, which could be easily extended to the case with different numbers of pieces. With $\mathcal{H}$ being the class of continuous PA functions with each max-affine component having $K$ pieces and parameters $\{\theta_{\iota}\}_{\iota\in[d]}$ satisfying $\left\lVert\theta_{\iota}\right\rVert_{\infty}\leq\mu$ for $\iota\in[d]$, the ERM problem (3) is formulated as the following parameter estimation problem
$\min_{\left\lVert\theta_{\iota}\right\rVert_{\infty}\leq\mu,\forall\iota\in[d]}\quad\frac{1}{n}\sum_{s=1}^{n}\varphi\left(\Big{(}\max_{k\in[K]}\left\{(\alpha^{k}_{\iota})^{\top}x^{s}+a^{k}_{\iota}\right\}-\max_{k\in[K]}\left\{(\beta^{k}_{\iota})^{\top}x^{s}+b^{k}_{\iota}\right\}\Big{)}_{\iota=1}^{d};y^{s}\right).$ (5)
In response to the aforementioned question, we intend to tackle two critical issues of the PADR-based ERM model in this paper: the consistency issue of the PADR-based ERM problem (5) and the computational issue for solving this large-scale, nonconvex, and nondifferentiable composite optimization problem. We address the first issue by quantifying the excess risk bound of our model, a standard procedure for model selection (Bartlett et al. 2002) in machine learning, and the second issue by developing a sampling-based algorithm with convergence guarantees, through which we demonstrate the ability of our proposed PADR-based ERM approach to effectively solve a broad class of SP problems with covariate information.
@@ 1.2 Literature Review
We present a literature review of data-driven methods for SP with covariate information and algorithmic advances in solving nonconvex optimization problems.
1.2.1 SP with Covariate Information.
@@
There are two main schemes of data-driven approaches in the literature of SP dealing with the latent dependency on covariate information: the sequential and integrated schemes. In the sequential scheme, also referred to as the predict-then-optimize (PO) scheme (Deng and Sen 2022, Kannan et al. 2022), one first predicts uncertain parameters and then solves the optimization problem composite with the prediction model. Several works such as Kao et al. (2009), Donti et al. (2017), and Elmachtoub and Grigas (2021) propose modifications of the loss function by accounting for the cost error of the downstream optimization problem. Since those modified loss functions involve optimal solutions to the decision-making problems that are parameterized by suspicious prediction models, such frameworks usually lead to bi-level nonconvex optimization problems and thus encounter substantial computational challenges in dealing with general constrained nonlinear programs.
@@
Approaches in the integrated scheme yield decisions via one-step data-driven optimization. Bertsimas and Kallus (2019) and Kallus and Mao (2022) approximate the conditional expected cost in (1) by a reweighted empirical cost. However, as claimed by Bertsimas and Koduri (2021), this framework is highly affected by the data with features close to the current observation and thus suffers from the curse of dimensionality. The DR-based ERM framework presented in (3) is another one-step approach in the integrated scheme. Due to its functional approximation in nature, this approach is relieved from the aforementioned issue.
Besides the classical LDR, several generalized LDRs are proposed for standard SP problems which lift random vectors into a higher-dimensional probability space to describe nonlinear relations (Chen et al. 2007, Bampou and Kuhn 2011, Georghiou et al. 2015). However, they tend to have unsatisfactory performance without cautious prior designs especially when the random vector is high-dimensional; moreover, except for the posterior estimation provided in Kuhn et al. (2011), their approximation accuracy lacks theoretical guarantees. The PADR with the parametric formula (4) employed in Bertsimas and Georghiou (2015) and this paper requires no prior designs of breakpoints as in Georghiou et al. (2015); instead, the breakpoints are implicitly determined by parameters in (4). Unlike the mixed-integer formulation in Bertsimas and Georghiou (2015) for multi-stage adaptive optimization problems, we treat the PADR-based ERM problem (5) as a special class of nonconvex and nondifferentiable optimization problems and exploit algorithmic advances to accommodate the large sample size and high-dimensional feature space.
@@
Several works on DR-based ERM methods are most relevant to this paper. Ban and Rudin (2018) employ LDR for the Newsvendor problem with nonasymptotic optimality guarantees under the assumption of the linear demand model. Notz and Pibernik (2021) and Bertsimas and Koduri (2021) study decision rules in reproducing kernel Hilbert spaces (RKHS), which we refer to as RKHS-DR, and provide asymptotic optimality guarantees for convex SP problems. Compared to the above references, our PADR-based ERM model is suitable for a much broader class of SP problems that can be nonconvex, and by following the statistical learning theory (Vapnik 1999, Mohri et al. 2018), we establish the nonasymptotic consistency result of our model without any parametric assumptions on the optimal decision rule or the latent dependency of the random variable $Y$ on the feature $X$.
1.2.2 Difference-of-convex (DC) and Composite DC Programming.
Due to the finite-sum composite structure of the objective in (5) with potentially large sample size $n$ and feature dimension $p$, it is prohibitively challenging to obtain a global minimizer to such a nonconvex and nondifferentiable optimization problem. Thus, our goal is to compute a promising candidate for a local minimizer via a sampling-based algorithm.
@@
When the cost function $\varphi(\cdot;y)$ is piecewise affine (e.g., the Newsvendor cost), the overall composite objective of (5) is also piecewise affine, which is a special DC function. Deterministic DC algorithms have been extensively studied for solving DC programs over the last several decades (Le Thi and Pham Dinh 2018). When the concave part of the DC decomposition is nondifferentiable, the basic DC algorithm converges to a critical point that satisfies the necessary condition for local optimality. To obtain a d(irectional)-stationary point that is the sharpest first-order stationary point for general nonconvex problems, Pang et al. (2016) study a class of nondifferentiable DC programs whose concave part is the maximum of a finite number of smooth convex functions and propose an enhanced DC algorithm which iteratively solves convex subproblems constructed by linearizing the $\varepsilon$-active piece of the concave part. Though sampling-based algorithms for stochastic DC programs have been recently studied in Nitanda and Suzuki (2017), Xu et al. (2019), and Le Thi et al. (2022) addressing smooth or partially smooth objectives, there still lacks a comprehensive study of the enhanced surrogation technique in stochastic DC algorithms to seek d-stationary solutions to the stochastic DC programs with nondifferentiable concave parts.
With a more general cost function that is not necessarily piecewise affine, the finite-sum composite structure of the objective lies at the cornerstone of algorithmic development for solving the PADR-based ERM problem (5). Cui et al. (2018) study the ERM problem with a convex composite DC structure and develop a majorization-minimization (MM) algorithm with an enhanced surrogation technique similar to Pang et al. (2016); however, such a deterministic algorithm requires iteratively solving finite-sum convex subproblems over all the data, raising concerns about its computational efficiency for solving large-scale ERM problems. The composite DC problem (5) can be regarded as a special class of compound SP studied in Liu et al. (2022) with finite support and solved by the stochastic majorization-minimization (SMM) algorithm proposed by the authors. Nonetheless, without the enhanced surrogation technique, SMM can only asymptotically compute a compound critical point when the inner concave function is nondifferentiable. Finally, we refer readers to Cui and Pang (2021) for a comprehensive study of modern nonconvex and nondifferentiable optimization, especially Chapters 6 and 7 for the study of stationary concepts and surrogation-based algorithms for nonconvex composite optimization problems.

### v4

@@
Focusing on stochastic programming (SP) with covariate information, this paper proposes an empirical risk minimization (ERM) method embedded within a nonconvex piecewise affine decision rule (PADR), which aims to learn the direct mapping from features to optimal decisions. We establish the nonasymptotic consistency result of our PADR-based ERM model for unconstrained problems, which illustrates the role of piece number in balancing the trade-off between the approximation and estimation errors. To solve the nonconvex and nondifferentiable ERM problem, we develop an enhanced stochastic majorization-minimization (ESMM) algorithm and establish the convergence to (composite strong) directional stationarity along with convergence rate analysis. We show that the proposed PADR-based ERM method applies to a broad class of nonconvex SP problems with theoretical consistency guarantees and computational traceability. Our numerical study demonstrates the effectiveness of PADR-based ERM methods with comparisons to state-of-the-art approaches under various settings, with significantly lower costs, less computation time, and robustness to feature dimensions and nonlinearity of the underlying dependency.
@@ 1 Introduction
With the fast-growing real-world applications and availability of large datasets, the field of stochastic programming (SP) has flourished with many theoretical and algorithmic advances (Birge and Louveaux 2011, Shapiro et al. 2021), in which the empirical probability distribution based on historical scenarios plays a central role in approximating the true distribution. However, it is common in practice that some covariate information (also referred to as contextual information or features) that may affect the random variable could be observed before the decision is made. In such situations, the empirical probability distribution and classical SP methodologies ignoring the covariate information leads to the poor performance. Given the covariate information $X=x\in\mathbb{R}^{p}$, an unconstrained SP is formulated as follows,
$\inf_{z\in\mathbb{R}^{d}}\quad\mathbb{E}_{Y}\left[\varphi({z},{Y})\mid X=x\right],$ (1)
where $z\in\mathbb{R}^{d}$ is a decision vector, $\varphi:\mathbb{R}^{d}\times\mathcal{Y}\to\mathbb{R}$ is a cost function, and $Y$ is a random variable supported by $\mathcal{Y}$ whose distribution is related to a feature vector $X$ supported by $\mathcal{X}\subseteq\mathbb{R}^{p}$. A simple instance is an inventory planning problem, in which $z$ represents order decisions for $d$ products, $\varphi$ represents the inventory cost, $Y$ represents uncertain demands, and $X$ represents covariate information such as seasonality, promotion, weather, past orders, etc. With historical scenarios of $(X,Y)$, a classical sample average approximation (SAA) model disregards the dependency of $Y$ on $X$ and thus may lead to inconsistent decisions; such deficiency is also observed in numerical results of Ban and Rudin (2018) and Bertsimas and Koduri (2021). Motivated by the latent dependency of the optimal solution to (1) on the feature vector, this paper aims to construct a data-driven decision rule that directly maps features to optimal decisions so as to approximately solve the SP problem (1) for any possible observation of the feature.
@@ 1.1 Problem setting
$\inf_{f\in\mathcal{F}}\quad R(f)\coloneqq\mathbb{E}_{X,Y}\left[\varphi({f(X)},{Y})\right],$ (2)
@@
where $\mathcal{F}$ contains all measurable functions mapping from the feature space $\mathcal{X}$ to the decision space $\mathbb{R}^{d}$. By the interchangeability principle (Shapiro et al. 2021, Theorem 7.80), if a mapping $f^{*}\in\mathcal{F}$ is optimal to problem (2), then the decision $f^{*}(x)$ is optimal to (1) for almost every $x\in\mathcal{X}$. In general, problem (2) is intractable due to the infinite-dimensional function space and the unknown joint distribution of $X$ and $Y$. It is common to consider an approximation method which constructs an empirical risk minimization (ERM) problem based on a dataset $\Xi\coloneqq\{(x^{s},y^{s})\}_{s=1}^{n}$ with the decision rule $f$ restricted to a parametric DR class $\mathcal{H}$ (also known as hypothesis class):
$\inf_{f\in\mathcal{H}}\quad\widehat{R}_{\Xi}(f)\coloneqq\frac{1}{n}\sum_{s=1}^{n}\varphi({f(x^{s})},{y^{s}}).$ (3)
@@
In this paper, we adopt the above DR-based ERM framework, recognizing that the efficacy of such a framework is influenced by the trade-off between approximation accuracy of $\mathcal{H}$ to the optimal DR and computational complexity of solving (3). For example, problem (3) with a convex cost function and linear decision rules (LDR) can be solved efficiently (Beutel and Minner 2012, Ban and Rudin 2018), while the LDR has difficulty describing the potential nonlinear relation between decisions and features. On the other side, a DR class consisting of neural networks could have a strong approximation capability (Qi et al. 2022); nevertheless, the resulting nonconvex problem (3) with a multi-layer composite structure is difficult to solve with theoretical optimality guarantees, and the model may face overfitting issues when the dataset is not large enough.
@@
Therefore, this paper aims to address a critical question: what could be an amenable hypothesis class for the DR-based ERM framework so as to maintain both high reliability and computational efficiency for solving SP with covariate information? To this end, we employ a piecewise affine decision rule (PADR) in the ERM problem (3) that is broader than LDR but simpler than many general nonlinear DRs. According to Scholtes (2012), any continuous piecewise affine (PA) function $f:\mathcal{X}\rightarrow\mathbb{R}$ can be expressed as the difference of max-affine functions:
$f(x,\theta)=\max_{k\in[K]}\left\{(\alpha^{k})^{\top}x+a^{k}\right\}-\max_{k\in[K]}\left\{(\beta^{k})^{\top}x+b^{k}\right\},$ (4)
@@
where $\theta\coloneqq\left(\alpha^{k},a^{k},\beta^{k},b^{k}\right)_{k=1}^{K}\in\mathbb{R}^{2K(p+1)}$ and $[K]\coloneqq\{1,\dots,K\}$. The PADR of such a form is beneficial because it requires minimal parameter tuning in multi-dimensional cases without the need to specify breakpoints for piecewise linear parts in each dimension, and it can asymptotically approximate any continuous functions according to Royset (2020). Throughout the theoretical analysis in this paper, we assume that the numbers of pieces in the two max-affine components of (4) are equal to simplify the exposition, which could be easily extended to the case with different numbers of pieces.
To address the aforementioned question, we intend to tackle two critical issues of the PADR-based ERM model for solving the unconstrained decision-making problem (1) in this paper: the consistency guarantees of the PADR-based ERM problem and the computational method for efficiently solving a large-scale, nonconvex, and nondifferentiable composite optimization problem. It is more challenging to extend the data-driven decision rule methods for constrained decision-making problems. Motivated by the scenario approximation method in Calafiore and Campi (2005), a straightforward approach is to impose the strengthened feasibility requirement over the observed scenarios in learning the data-driven decision rule. In Appendix 9, we illustrate that the obtained decision rule at a new scenario is feasible with high probability and is asymptotically consistent with the original SP problem, and we also discuss several special cases where the feasibility of the decision rule can be guaranteed.
In the rest of this section, we give a brief literature review and present the distinguished features of the proposed PADR method over other relative work. We will also summarize major contributions of the present paper.
@@ 1.2 Literature review
There are two main schemes of data-driven approaches in the literature of SP dealing with the latent dependency on covariate information: the sequential and integrated schemes. In the sequential scheme, also referred to as the predict-then-optimize (PO) scheme (Deng and Sen 2022, Kannan et al. 2022), one first predicts uncertain parameters and then solves the optimization problem composite with the prediction model. Several works such as Kao et al. (2009), Donti et al. (2017), Elmachtoub and Grigas (2021), and Qi et al. (2021a) propose modifications of the loss function by considering the cost error of the downstream optimization problem. Since those modified loss functions involve optimal solutions to the decision-making problems that are parameterized by prediction models, such frameworks usually lead to bi-level nonconvex optimization problems and thus encounter substantial computational challenges in dealing with nonlinear constrained programs.
Approaches in the integrated scheme yield decisions via one-step data-driven optimization. Prescriptive methods by Bertsimas and Kallus (2019) and Kallus and Mao (2023) approximate the conditional expected cost in (1) through reweighing empirical costs. Elmachtoub et al. (2023) point out that, when the model class is well-specified and there is sufficient data, methods that directly estimate the conditional distribution outperform those that consider the downstream optimization problem; in a misspecified scenario, the reverse holds.
@@
However, as claimed by Bertsimas and Koduri (2021), the prescriptive methods are highly affected by the data with features close to the current observation and thus suffers from the curse of dimensionality. The DR-based ERM framework presented in (3), which is another one-step approach in the integrated scheme, is relieved from the aforementioned issue due to its functional approximation in nature. For constrained SP with general nonlinear constraints, the current study of decision rule methods could only guarantee feasibility with high probability, and the predict-then-optimize framework may be more suitable for solving the contextual problems when the feasibility is essential to decision-makers.
@@
Several works on DR-based ERM methods are most relevant to this paper. Ban and Rudin (2018) employ LDR for the Newsvendor problem with nonasymptotic optimality guarantees under the assumption of the linear demand model. Notz and Pibernik (2021) and Bertsimas and Koduri (2021) study decision rules in reproducing kernel Hilbert spaces (RKHS), which we refer to as RKHS-DR, and provide asymptotic optimality guarantees for convex SP problems. Compared to the above existing DR methods, given the asymptotic approximation of PA functions to any continuous functions, the PADR method is applicable to more general decision-making problems with nonconvex cost function or nonlinear dependency with asymptotic optimality.
Besides the classical LDR, several generalized LDRs are proposed for standard SP problems which lift random vectors into a higher-dimensional probability space to describe nonlinear relations (Chen et al. 2007, Bampou and Kuhn 2011, Georghiou et al. 2015). However, they tend to have unsatisfactory performance without cautious prior designs especially when the random vector is high-dimensional; moreover, they usually lack theoretical guarantees on approximation accuracy except for the posterior estimation provided in Kuhn et al. (2011).
Regarding solving the PADR-based ERM problem with potentially large sample size $n$, it is prohibitively challenging to obtain a global minimizer to such a coupled nonconvex and nondifferentiable optimization problem. As provided by Bertsimas and Georghiou (2015, Section 2.4), the PADR-based ERM problem can be formulated as a mixed-integer nonlinear program using Big-M method with additional $2ndK$ binary variables to describe the PADR of interest. For a PADR-based ERM problem of the modest size with data size $n=1000$, the obtained mixed-integer nonlinear program could have $10^{4}\sim 10^{5}$ binary variables and the state-of-the-art solvers are hard to solve within a few hours to achieve an approximately optimal solution.
Instead of seeking global optimality, majorization-minimization (MM) methods provide a unified framework that is practically efficient and theoretically convergent for solving the PADR-based ERM problem with the finite-sum composite objective. MM methods iteratively construct convex subproblems with an upper surrogate objective function for obtaining “stationary” solutions. According to variational analysis by Rockafellar and Wets (2009), there are many kinds of stationary concepts which differ in the ability to characterize necessary optimality conditions for local/global optimality. As pointed in Cui and Pang (2021, Chapter 7), MM algorithms may converge to different types of stationary solutions depending on the properties of surrogate and objective functions. Algorithms with convergence to a stronger stationarity is favorable not only theoretically but also computationally because the obtained solution from numerical implementation with random initialization has the greater potential to be a global optimum.
@@
Difference-of-convex (DC) algorithms, which is a special MM method, have been extensively studied over the last several decades (Le Thi and Pham Dinh 2018) and can compute a d(irectional)-stationary point for DC programs with a differentiable concave component. When the cost function $\varphi(\cdot,y)$ is convex or piecewise affine, the PADR-based ERM problem can be formulated as a nondifferentiable DC program for which DC algorithms including Tao and An (1997) and Lipp and Boyd (2016) are applicable. However, due to the nondifferentiable concave component, DC algorithms for the ERM problem converge to a critical point whose defining condition is dependent on the dc decomposition form and is weaker than d(irectional)-stationarity. For nondifferentiable DC or composite DC problems, Pang et al. (2016) and Cui et al. (2018) propose an enhanced surrogation technique that iteratively solves convex subproblems that are constructed based on $\varepsilon$-active pieces of the concave component. However, the enhanced MM algorithm for large-scale ERM problems requires iteratively solving finite-sum convex subproblems over all the data, which raises the fundamental issue of computational efficiency. Sampling-based algorithms for stochastic DC programs have been recently studied in Nitanda and Suzuki (2017), Xu et al. (2019), Le Thi et al. (2022) addressing smooth or partially smooth objectives, and in Liu et al. (2022) addressing nonconvex and nonsmooth compound SP to compute critical solutions. There still lacks a comprehensive study on whether the $\varepsilon$-active set technique is amenable for sampling-based methods to enhance the convergence properties for nondifferentiable stochastic DC programs and PADR-based ERM problems.